Solve ( x2 – 2x )( x + 3) = -2x ( x + 1)
Answer by sowmya(32)

( x2 – 2x )( x + 3) = -2x ( x + 1)
x(x-2)(x+3)=-2x(x+1)
x(x-2)(x+3)+2x(x+1)=0
x{(x-2)(x+3)+2(x+1)}=0
x=0
or
(x-2)(x+3)+2(x+1) =0
x^2+x-6+2x+2 = 0
x^2+3x-4=0
x^2+4x-x-4=0
(x-1)(x+4)=0
Hence the solutions are
x=0, x=1 or x=-4